4 - 2(x-5) + 4x = 8
I keep getting -1 but my teacher said it was wrong.

Answer by ankor@dixie-net.com(22740)

4 - 2(x-5) + 4x = 8
multiply the terms inside the brackets, it's -2, therefore you change the signs
4 - 2x + 10 + 4x = 8
Combine like terms
-2x + 4x + 4 + 10 = 8
2x + 14 = 8
Subtract 14 from both sides
2x = 8 - 14
2x = -6
x = -6/2
x = -3
:
:
Check this in the original equation
4 - 2(-3-5) + 4(-3) = 8
4 - 2(-8) - 12 = 8
4 + 16 - 12 = 8
8 = 8; confirms our solution